Circular No. 6643

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970228 D. W. Fox, J. M. Kommers, and W. H. G. Lewin, Massachusetts In-

stitute of Technology; and E. A. Magnier, University of Washington, report: “We have analyzed the Hubble Space Telescope (HST) Planetary Camera images (Sahu et al., IAUC 6606, 6619), taken with the F606W (wide V) and F814W (I) filters, of the optical transient associated with GRB 970228 (IAUC 6584) in an attempt to characterize the extended emission first reported on IAUC 6588 and later found to have faded below detection level (IAUC 6631). We find that the profile of the transient is well characterized by a synthetic HST point-spread function (see

http://scivax.stsci.edu/~krist/tinytim.html); subtracting the fit-

ted contribution of the optical transient, and excluding the region within 3 pixel lengths of its peak, we performed aperture photometry on an ellipse containing the bulk of the apparent emission, with semimajor axis 0′′.73 and semiminor axis 0′′.55 (16×12 pixels), centered at α = 5^{h}01^{m}46^{s}.65, δ = +11^{o}46′54′′.0 (equinox 2000.0), and having major axis oriented 16^{o} west of north. The sky background is determined from a Gaussian fit to a histogram of pixel intensities from the central 5′′.8 × 5′′.8 of each image. The extended source is detected at > 3σ significance in all but one image (Mar. 26.5 UT, F814W); the excess countrates (in DN/s; see Holtzman et al. 1995, PASP 107, 1065, for conversion to physical units) that we associate with it are as follows: 0.078 ± 0.015 (Mar. 26.4, F606W), 0.027 ± 0.019 (Mar. 26.5, F814W), 0.074 ± 0.016 (Apr. 7.2, F606W), 0.064 ± 0.019 (Apr. 7.3, F814W). We thus find no evidence for a change in the intensity or F606W– F814W color of the extended emission between the two HST observations. Assuming a normal distribution of our errors, we find that any change in the intensity of the extended emission over this time interval lies (with 90percent confidence) between −0.48 and +0.60 mag in F606W, and between −2.42 and +0.40 mag in F814W.”

SUPERNOVA 1997by IN ANONYMOUS GALAXY R. Kirshner and S. Jha, Center for Astrophysics, report that a spec-

trum of SN 1997by, taken by P. Berlind at the Whipple Observatory’s 1.5-m telescope on Apr. 30.28 UT, confirms that it is a type-Ia supernova near maximum at a redshift of z = 0.045 (cf. IAUC 6640). This matches well the Abell Cluster 1736 redshift of z = 0.046.

1997 May 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green